# Mitigating Cross-Modal Distraction and Ensuring Geometric Feasibility via Affordance-Guided, Self-Consistent MLLMs for Food Preparation Task Planning

### v1

@@ Abstract
We study Multimodal Large Language Models (MLLMs) with in-context learning for food preparation task planning. In this context, we identify two key challenges: cross-modal distraction and geometric feasibility. Cross-modal distraction occurs when the inclusion of visual input degrades the reasoning performance of a MLLM. Geometric feasibility refers to the ability of MLLMs to ensure that the selected skills are physically executable in the environment. To address these issues, we adapt Chain of Thought (CoT) with Self-Consistency to mitigate reasoning loss from cross-modal distractions and use affordance predictor as skill preconditions to guide MLLM on geometric feasibility. We construct a dataset to evaluate the ability of MLLMs on quantity estimation, reachability analysis, relative positioning and collision avoidance. We conducted a detailed evaluation to identify issues among different baselines and analyze the reasons for improvement, providing insights into each approach. Our method reaches a success rate of 76.7% on the entire dataset, showing a substantial improvement over the CoT baseline at 36.7%.
@@ I INTRODUCTION
Imagine walking into a high-end tofu pudding shop where a service robot is tasked with preparing customized dessert orders. A table in front of the robot is set with bowls of tofu pudding and various toppings. Customers interact with the robot using natural language, specifying their preferences, such as ‘Place some mung beans onto my tofu pudding.’ To fulfill these requests, the robot must accurately interpret human instructions, ground the specified food items, and perform the corresponding manipulation skills to complete tofu pudding preparation.
Interpreting human instructions and converting them into the corresponding action sequences remains a challenge in robotics. Task planning [3, 4, 5, 6, 7, 8] decomposes a given goal specification into subtasks based on the available robot skill set. To handle goal specifications expressed in natural language, Large Language Models (LLMs) have been investigated for their ability to generate structured plans from textual input. More recently, Multimodal Large Language Models (MLLMs) have extended LLMs by incorporating additional modalities, such as visual inputs, enabling them to perceive and interpret the world.
In this work, we study MLLMs with in-context learning for food preparation task planning. Due to the difficulties in recovering from spillage in real-world settings and the lack of datasets in simulation that resemble our food preparation scenario, we construct our dataset in a simulated environment to evaluate the abilities of MLLMs. Specifically, we choose IsaacGym [9] as our simulator due to its high-fidelity in rendering a wide range of food items.
We identify cross-modal distraction and geometric feasibility as the main challenges to applying MLLMs for food preparation task planning. Reasoning across different modalities has been an active area of research in the community  [10, 11]. However, challenges such as visual hallucination and illusion highlight that incorporating image inputs can sometimes produce unintended or misleading results [12, 13]. Our experiments reveal that for tasks not requiring visual input, adding images to the same prompt distracts the MLLM, degrading its reasoning performance. We term this phenomenon cross-modal distraction. For example, in tasks that require only semantic reasoning, incorporating an image observation causes MLLMs to generate infeasible skill sequences, overlook previous actions, or scoop twice despite being instructed to scoop only once.
Additionally, we find that MLLMs struggle with reasoning about geometric feasibility, which involves ensuring that the planned sequence of robot actions is physically executable within the spatial constraints of the environment. For example, in Figure LABEL:fig-teaser, MLLM fails to recognize that there is a bowl in front of the dumbwaiter door and attempts to open the dumbwaiter directly.
A natural question arises: how can we address the limitations of MLLMs in food preparation task planning within an in-context learning setting?
@@
Our approach uses Chain of Thought with self-consistency to counter performance drops from cross-modal distractions and leverages skill affordance to provide geometric feasibility information for MLLM. We study task planning under closed-loop setting, where the robot iteratively selects actions using the new image input. Chain of Thought prompting [20] helps MLLM to reason through problems step by step, leading to better responses. Self-Consistency Verification in MLLM checks whether the model generates reliable outputs across multiple runs, enabling it to detect anomalies and maintain stable performance. Skill affordance refers to the likelihood that a skill can be successfully executed given a specific environment state and acts as the action precondition for executing certain skills. We encode crucial information for identifying geometric feasibility as preconditions, enabling verification of whether the decisions made by MLLM are feasible. Our contributions are as follows:
@@
    We construct a dataset designed for task planning in food preparation, categorizing different scenarios to evaluate various robotic capabilities.
@@
    We adapt CoT with Self-Consistency to closed-loop task planning settings to combat reasoning loss caused by cross-modal distractions
@@
    We build up action preconditions with predicates to enable MLLM to take geometric feasibility into consideration.
@@ II-B Reasoning Ability of Large Models
Self-Consistency methods employ a ”sample and select” approach, where large models generate several possible answers and determine the final output with majority vote [24, 25]. We extend this concept to closed-loop task planning in robotics, using the output of previous iterations as samples to assess the LLM’s consistency at the current iteration.
@@ III-B Pipeline Overview
Our planning pipeline (Fig. 2) consists of three main stages: 1) MLLM Planning, 2) Self-Consistency Verification, and 3) Skill Affordance and Replanning. In the first stage, we leverage Chain of Thought (CoT) [23] prompting to enhance the reasoning ability of MLLM. The latter two stages address the key challenges: cross-modal distraction and geometric feasibility. Self-Consistency Verification is designed to mitigate the degraded reasoning ability and stability caused by cross-modal distraction. Skill Affordance and Replanning focus on evaluating geometric feasibility, preventing infeasible skill executions through precondition checks, and recovering from infeasibility by utilizing structured feedback.
@@ III-C Zero-Shot CoT
In the first stage of our planning pipeline, we provide environmental information and examples to help MLLMs understand the input-output format and objective. To enhance reasoning ability, we integrate a zero-shot CoT [23] approach, prompting the MLLM to first generate an explanation and description of the sub-goal, reason through the current iteration, and then determine the final skill sequence decision. Here is an example response demonstrating how our approach operates: {mdframed}[linewidth=1pt] Description: The blue bowl appears to be fuller, so I will scoop from the blue bowl and drop the contents into the pink bowl.
@@ III-D Self-Consistency Verification
Inspired by previous works [24, 25], we design a mechanism to evaluate the consistency of MLLM across iterations. We instruct MLLM to produce a full sequence of skills instead of a single skill. The skill selected for the current iteration will be considered for execution, while the remaining skills serve as the source of consistency checks and are stored into a sequence buffer when the skill selected for execution is being executed. At each iteration, the selected skill for execution is checked against the majority of previous decisions. If discrepancy exists, MLLM is prompted to reconsider and choose between the conflicting skills.
For example, MLLM generates a sequence $s_{t}$ with $\pi_{t}$ as scoop at iteration $T=t$. However, the most frequently chosen skill at $T=t$ from the previous sequence buffer is move_to_white_bowl. Due to this discrepancy, the MLLM must resolve the conflict and select a final decision. If the MLLM chooses move_to_white_bowl and it is validated by the Skill Affordance module, robot will execute this skill, and the suquence buffer will be updated with $s_{t}$.
@@ III-E Skill Affordance and Replan
To enable the MLLM to assess the skill affordance, we design predicates [8], which are binary values that answer specific questions about the environment. These predicates are combined into preconditions to evaluate the affordance of each skill. Here is the set of designed predicates along with their corresponding descriptions:
@@ IV EXPERIMENT
IV-A Dataset
[…]
[width=0.8]img/experiment.png
TABLE I: Experiment results. SC stands for Self-Consistency and SA stands for Skill Affordance.
We construct a dataset in tabletop food-serving scenario in IsaacGym [9]. The dataset consists of five categories of tasks, each containing 30 unique configurations, resulting in a total of 150 configurations. These configurations vary in instruction, container position and color, as well as the type and amount of food inside the container. In addition to the four categories designed to pose different challenges to MLLM, we create another category to evaluate its reasoning ability solely within the semantic domain. The following are descriptions of the five categories:
@@ IV-A Dataset
The environment includes a Franka Emika Panda 7-DoF robot arm, a table, a tool holder, and a dumbwaiter. Several bowls are placed on the table, some containing tofu pudding and others containing some toppings. There are seven types of toppings that differ in color.
@@ IV-B Baselines
To evaluate our approach, we compare against several baselines, as outlined below.
  • Naive LLM: To evaluate the impact of cross-modal input on reasoning ability, we test the ability of MLLM without image input on our dataset.
@@
  • CoT: We apply Chain of Thought prompting on the Naive MLLM planner by asking the MLLM to describe what it should achieve in the current iteration before selecting a skill.
  • CoT+SC: We implement the Self-Consistency mechanism on the CoT baseline.
@@ V RESULT AND DISCUSSION
Table I shows the success rate of each method and dataset categories. Our approach achieve an overall success rate of 76.7%. The result shows that 1) CoT with Self-Consistency enhances MLLM’s reasoning ability, resulting in improved performance on tasks in Semantic Reasoning, Quantity Estimation, and Relative Positioning. 2) Skill affordance enable robots to understand geometric feasibility, allowing them to assess action preconditions and successfully complete tasks in Reachability Analysis and Collision Avoidance.
@@ VI CONCLUSIONS
In this work, we explore the application of Multimodal Large Language Models (MLLMs) for task planning in food preparation, identifying key challenges associated with cross-modal distraction and geometric feasibility. Our findings highlight that while MLLMs demonstrate strong reasoning abilities, their performance can degrade when unnecessary visual input introduces distractions. To mitigate this, we applied Chain of Thought reasoning with Self-Consistency, improving the model’s robustness against cross-modal interference. Additionally, we integrated skill affordance as a physics engine, enabling the model to assess geometric feasibility when planning robotic actions.
Through the construction of a dedicated dataset and evaluation in a closed-loop task planning setting, we demonstrated the effectiveness of our approach in handling cross-modal distraction and recognizing geometric feasibility in food preparation tasks. Our contributions include dataset development, the adaptation of reasoning techniques for multimodal task planning, and the incorporation of action preconditions to enhance feasibility assessments.
@@ B Prompt Example
System prompt: A fixed and structured system prompt provided to the MLLM, organized into six sections and including two examples for in-context learning. {mdframed}[linewidth=1pt] # Scenario
@@ 3. Scooping limitations
Example 1: {adjustwidth}1.5em0pt Skill set: [’scoop’, ’stir’, ’drop food’, ’pull bowl closer’, ’open dumbwaiter’, ’close dumbwaiter’, ’start dumbwaiter’, ’put bowl into dumbwaiter’, ’DONE’, ’grasp spoon’, ’put spoon back’, ’move to white bowl’, ’move to green bowl’, ’move to blue bowl’] Initial object list: [’white bowl (with mung beans)’, ’green bowl (with mung beans)’, ’blue bowl (with tofu pudding)’]
@@
Example 2: {adjustwidth}1.5em0pt Skill set: [’scoop’, ’stir’, ’drop food’, ’pull bowl closer’, ’open dumbwaiter’, ’close dumbwaiter’, ’start dumbwaiter’, ’put bowl into dumbwaiter’, ’DONE’, ’grasp spoon’, ’put spoon back’, ’move to purple bowl’, ’move to red bowl’]
@@
- Iteration 2 to 5 demonstrates how to scoop from red bowl to purlple bowl once, and iteration 6 to 9 repeats it. Example user prompt: Includes environmental information, task instructions, and feedback messages from the Skill Affordance module, along with the current observation.
@@
{mdframed}
[linewidth=1pt] Your task: {adjustwidth}1.5em0pt Skill set: [’A. scoop’, ’B. stir’, ’C. drop food’, ’D. pull bowl closer’, ’E. open dumbwaiter’, ’F. close dumbwaiter’, ’G. start dumbwaiter’, ’H. put bowl into dumbwaiter’, ’I. DONE’, ’J. grasp spoon’, ’K. put spoon back’, ’L. move to white bowl’, ’M. move to blue bowl’]

### v2

# Mitigating Cross-Modal Distraction and Ensuring Geometric Feasibility via Affordance-Guided and Self-Consistent MLLMs for Task Planning in Instruction-Following Manipulation

@@ Abstract
We investigate the use of Multimodal Large Language Models (MLLMs) with in-context learning for closed-loop task planning in instruction-following manipulation. We identify four essential requirements for successful task planning: quantity estimation, reachability analysis, relative positioning, and collision avoidance. However, existing benchmarks fail to support holistic evaluation across all these aspects. To address this gap, we introduce QuARC (Quantity, Analysis, Relative positioning, Collision), a new benchmark based on a food preparation scenario that integrates all four challenges. Using QuARC, we reveal two major limitations of current MLLMs: cross-modal distraction and geometric infeasibility. To tackle these, we adapt Chain-of-Thought with Self-Consistency to mitigate reasoning loss from cross-modal distractions and incorporate an affordance predictor to guide planning based on geometric feasibility. Our comprehensive evaluation analyzes performance across multiple baselines and explains sources of improvement. Our method achieves a 76.7% success rate on the benchmark, significantly outperforming the ViLa baseline (36.7%), without requiring additional finetuning. Code and dataset are available at https://hcis-lab.github.io/Affordance-Guided-Self-Consistent-MLLM
[…]
! Benchmark Type Quantity Estimation Reachability Analysis Relative Positioning Collision Avoidance ALFRED [4] Mobile Manipulation BEHAVIOR-1K [5] Mobile Manipulation CALVIN [6] Manipulation TableEnv [12] Manipulation QuARC (Ours) Manipulation : requirement supported  ✗: not supported
TABLE I: Comparison of existing benchmarks across four essential requirements, i.e., quantity estimation, reachability analysis, relative positioning, and collision avoidance.
@@ I INTRODUCTION
Imagine entering a tofu pudding shop where an embodied agent prepares customized dessert orders. In front of the agent is a table filled with bowls of tofu pudding and various toppings. Customers interact using natural language, expressing preferences like, “place a scoop of toppings from the fuller bowl into the yellow bowl.” To complete this long-horizon task, the agent must interpret the instruction, ground the specified food items, execute manipulation skills, evaluate action feasibility, and adapt its behavior accordingly. This demands a holistic understanding and reasoning, posing a significant challenge for instruction-following embodied AI [2, 3].
We identify four essential requirements for successful task execution: quantity estimation, reachability analysis, relative positioning, and collision avoidance. Given the inherent ambiguity of natural language, quantity estimation and relative positioning become particularly difficult when instructions are vague. For example, instructions “scoop from the fuller bowl” or “place the food into the bowl next to the tool holder” involve ambiguity. In such cases, the agent must semantically reason the instruction, perceive the environment, and ground referred objects across visual and linguistic modalities. Additionally, reachability analysis ensures that the agent can position its end-effector within kinematic limits to interact with the target, while collision avoidance guarantees that planned motions do not cause unintended contact with surrounding objects.
To systematically study and evaluate these essential requirements, we propose QuARC (Quantity, Analysis, Relative positioning, Collision), a new benchmark entered on a food preparation use case that integrates all four aspects.
As summarized in Table I, existing benchmarks typically focus on isolated aspects and fall short of enabling comprehensive evaluation of instruction-following task planning under a closed-loop setting. We construct QuARC in simulation using IsaacGym [14] due to the difficulty of recovering from spillage in real-world settings and the lack of existing simulation benchmarks resembling our food preparation scenario. Moreover, this design choice enables reproducible and fair comparisons among different instruction-following manipulation.
Interpreting human instructions and translating them into actionable sequences remains a significant challenge. Recent advances in Large Language Models (LLMs) have demonstrated their potential to generate structured plans from natural language input. Building on this, Multimodal Large Language Models (MLLMs) incorporate additional modalities, such as visual input, enabling models to perceive and reason about the physical world. In this work, we explore the use of MLLMs with in-context learning for task planning in instruction-following manipulation. This approach leverages general-purpose models with minimal overhead, significantly reducing the cost, time, and effort required to adapt to new tasks and environments.
We identify Cross-Modal Distraction and Geometric Feasibility as the main challenges to applying MLLMs using QuARC. While reasoning across modalities is a well-studied area [15, 16], prior work has shown that incorporating visual inputs can sometimes lead to hallucinations or misleading outputs [17, 18]. Our experiments reveal that in tasks where visual input is unnecessary, including images in the prompt can distract the MLLM and degrade reasoning performance. This is an effect we term Cross-Modal Distraction. Even in tasks requiring only visual inputs, they can cause the model to produce infeasible action sequences, omit previous steps, or repeat actions, such as scooping twice when instructed to scoop once. This instability is especially problematic in closed-loop execution, where a single error can propagate, and in instruction-following manipulation, where mistakes may lead to irreversible outcomes.
Geometric feasibility is a longstanding challenge in robotics, especially in manipulation and task planning. It involves ensuring that a planned sequence of actions is physically executable within the environment’s spatial constraints. We evaluate MLLMs on two core aspects of ensuring geometric feasibility: (1) determining whether the target object is within the robot’s kinematic reach for successful interaction, and (2) avoiding collisions when generating action sequences. Our experiments reveal that MLLMs consistently struggle in both areas, often producing plans that violate spatial constraints, misidentify targets, or result in unintended collisions. For instance, as shown in Figure LABEL:fig-teaser, the model fails to recognize a bowl blocking the dumbwaiter door and attempts to open it directly, causing the bowl to spill. This raises a critical question: how can we address such limitations within an in-context learning setting?
@@
To this end, we propose to a new MLLMs with in-context learning model based on Chain of Thought with self-consistency to counter performance drops from cross-modal distractions and leverages skill affordance to provide geometric feasibility information for MLLM. We study task planning under closed-loop setting, where the robot iteratively selects actions using the new image input. Chain of Thought prompting [25] helps MLLM to reason through problems step by step, leading to better responses. Self-consistency verification in MLLM checks whether the model generates reliable outputs across multiple runs, enabling it to detect anomalies and maintain stable performance. Skill affordance refers to the likelihood that a skill can be successfully executed given a specific environment state and acts as the action precondition for executing certain skills. We encode crucial information for identifying geometric feasibility as preconditions, enabling verification of whether the decisions made by MLLM are feasible.
Our contributions are as follows:
@@
    We introduce QuARC, a benchmark that captures the four core requirements for effective task planning in instruction-following manipulation
@@
    We identify cross-modal distraction and geometric infeasibility as key challenges when applying MLLMs to closed-loop instruction-following manipulation
@@
    We leverage CoT with Self-Consistency and incorporate action preconditions via predicate-based representations for closed-loop instruction-following manipulation
  4. 4.
    Extensive evaluations and ablation studies empirically validate the effectiveness of our proposed approach
@@ II-B Reasoning Ability of Large Models
Self-Consistency methods employ a ”sample and select” approach, where large models generate multiple reasoning chains on a question and determine the final output with majority vote [29, 30]. Instead of generating multiple reasoning paths at each iteration, we reuse the outputs from previous iterations as samples for Self-Consistency. This allows the system to stabilize predictions without requiring additional prompts or redundant sampling.
@@ III-B Pipeline Overview
Our planning pipeline (Fig. 2) consists of three main stages: (1) MLLM Planning, (2) Self-Consistency Verification, and (3) Skill Affordance and Replanning. In the first stage, we leverage Chain of Thought (CoT) [28] prompting to enhance the reasoning ability of MLLM. The latter two stages address the key challenges: cross-modal distraction and geometric feasibility. Self-Consistency Verification is designed to mitigate the degraded reasoning ability and stability caused by cross-modal distraction. Skill Affordance and Replanning focus on evaluating geometric feasibility, preventing infeasible skill executions through precondition checks, and recovering from infeasibility by utilizing structured feedback.
@@ III-C Zero-Shot CoT
In the first stage of our planning pipeline, we provide environmental information and examples to help MLLMs understand the input-output format and objective. To enhance reasoning ability, we integrate a zero-shot CoT [28] approach, prompting the MLLM to first generate an explanation and description of the sub-goal, reason through the current iteration, and then determine the final skill sequence decision. Here is an example response demonstrating how our approach operates: Description: The blue bowl appears to be fuller, so I will scoop from the blue bowl and drop the contents into the pink bowl.
@@ III-D Self-Consistency Verification
Self-Consistency methods [29, 30] focus on applying consistency techniques to a single reasoning problem. The most intuitive way to extend Self-Consistency to closed-loop task planning is to treat each question as an independent problem. This approach requires sampling multiple Chain-of-Thought (CoT) outputs at each step and selecting the final result via majority voting. While this may indeed help stabilize the predictions of MLLMs, it significantly increases the number of queries required, leading to higher computational cost.
We propose a mechanism to evaluate the consistency of MLLMs across iterative reasoning steps without introducing additional sampling. We instruct MLLM to produce a full sequence of skills instead of a single skill. The skill selected for the current iteration will be considered for execution, while the remaining skills serve as the source of consistency checks and are stored into a sequence buffer when the skill in current iteration is confirmed for execution. At each iteration, the selected skill for execution is checked against the majority of previous decisions. If discrepancy exists, MLLM is prompted to reconsider and choose between the conflicting skills.
For example, let $S=\{s_{i}\}$ denote the set of prior reasoning sequences, where $s_{i}^{n}$ represents the skill selected for step $n$ in the $i$-th iteration. At step $t$, MLLM generates a sequence $s_{t}$ where $s_{t}^{t}$ is scoop. However, the most frequently chosen skill
$\arg\max_{a}\sum_{i<t}\mathbb{I}[s_{i}^{t}=a]$
at time step $t$ is move_to_white_bowl. To resolve this discrepancy, MLLM is prompted to compare the two skill sequences and determine the final decision. The selected skill then serves as the input for the next stage of the pipeline. In closed-loop task planning, skills are selected iteratively to exploit newly observed environment information. Comparing the current iteration with the previous majority decision instead of relying on a simple majority vote allows the MLLM to determine whether the discrepancy is caused by model instability or genuine changes in the environment.
@@ III-E Skill Affordance and Replan
Predicates [13] are binary values that answer specific questions about the environment. We design predicates for our environment to enable the MLLM to assess the skill affordance. These predicates are combined into preconditions to evaluate the affordance of each skill. Here is our predicate design:
@@ IV EXPERIMENT
IV-A Benchmark
To systematically study and evaluate our approach, we propose QuARC (Quantity, Analysis, Relative positioning, Collision), a new benchmark for food preparation that assesses the instruction-following capabilities of robotic systems. QuARC is built on IsaacGym [14] and comprises five categories of tasks: Semantic Reasoning, Quantity Estimation (Qu), Reachability Analysis (A), Relative Positioning (R), and Collision Awareness (C). The Semantic Reasoning category evaluates reasoning ability purely within the semantic domain, which is a fundamental prerequisite before assessing performance in both semantic and visual domains. Each category contains 30 unique configurations, yielding a total of 150 configurations. These configurations vary in the provided instructions, container positions and colors, as well as the type and amount of food in each container. The following are descriptions of the five categories:
@@ IV-A Benchmark
The environment is designed to simulate a realistic tabletop food preparation scenario and includes a Franka Emika Panda 7-DoF robot arm mounted in front of a worktable. A tool holder is positioned on one side of the table to store utensils, while a dumbwaiter is placed adjacent to the workspace to enable ingredient delivery and removal. Several bowls are arranged on the table in varying configurations. Some bowls contain tofu pudding, while others contain a variety of toppings. In total, there are seven types of toppings, each distinguished by its color to facilitate visual recognition. The spatial arrangement, bowl colors, and topping types can vary across tasks, introducing diversity in both visual appearance and physical positioning, which in turn increases the complexity of perception and manipulation.
@@ IV-B Baselines
We compare against several baselines to evaluate our approach.
  • Naive LLM: To evaluate the impact of cross-modal input on reasoning ability, we test the ability of MLLM without image input on our benchmark. Note that it is expected to perform well only in Semantic Reasoning tasks, where the solutions can be derived purely from text.
@@
  • ViLa (CoT): According to ViLa [31], we apply CoT on Naive MLLM planner. We ask MLLM to describe what it should achieve in current iteration before selecting a skill.
  • ViLa+SC: We implement the Self-Consistency mechanism on the ViLa baseline.
@@ IV-C Setup and Implementation
[…]
[width=0.8]img/experiment.png
TABLE II: Experiment results. SC stands for Self-Consistency and SA stands for Skill Affordance.
@@ V RESULT AND DISCUSSION
Table II shows the success rate of each method and dataset categories. Our approach achieve an overall success rate of 76.7%. The result shows that 1) CoT with Self-Consistency enhances MLLM’s reasoning ability, resulting in improved performance on tasks in Semantic Reasoning, Quantity Estimation, and Relative Positioning. 2) Skill affordance enable robots to understand geometric feasibility, allowing them to assess action preconditions and successfully complete tasks in Reachability Analysis and Collision Avoidance.
@@ VI CONCLUSIONS
In this work, we explore the application of Multimodal Large Language Models (MLLMs) for task planning in instruction following manipulation, identifying four essential requirments associated with two challenges, cross-modal distraction and geometric feasibility. Our findings highlight that while MLLMs demonstrate strong reasoning abilities, their performance can degrade when unnecessary visual input introduces distractions. To mitigate this, we applied Chain of Thought reasoning with Self-Consistency, improving the model’s robustness against cross-modal interference. Additionally, we integrated skill affordance as a physics engine, enabling the model to assess geometric feasibility when planning robotic actions.
Through the construction of a dedicated benchmark and evaluation in a closed-loop task planning setting, we demonstrated the effectiveness of our approach in handling cross-modal distraction and recognizing geometric feasibility in food preparation tasks. Our contributions include benchmark development, the adaptation of reasoning techniques for multimodal task planning, and the incorporation of action preconditions to enhance feasibility assessments.
@@ VII LIMITATIONS and FUTURE WORK
ACKNOWLEDGMENT
This work was supported in part by the Center for Intelligent Team Robotics and Human-Robot Collaboration under the “Top Research Centers in Taiwan Key Fields Program” of the Ministry of Education (MOE), Taiwan, and in part by the National Science and Technology Council, Taiwan, under grants NSTC 113-2813-C-A49-019-E, 114-2218-E-011-003, and 114-2221-E-A49-009.
@@ B Prompt Example
System prompt: A fixed and structured system prompt provided to the MLLM, organized into six sections and including two examples for in-context learning. # Scenario
@@ 3. Scooping limitations
Example 1: Skill set: [’scoop’, ’stir’, ’drop food’, ’pull bowl closer’, ’open dumbwaiter’, ’close dumbwaiter’, ’start dumbwaiter’, ’put bowl into dumbwaiter’, ’DONE’, ’grasp spoon’, ’put spoon back’, ’move to white bowl’, ’move to green bowl’, ’move to blue bowl’] Initial object list: [’white bowl (with mung beans)’, ’green bowl (with mung beans)’, ’blue bowl (with tofu pudding)’]
@@
Example 2: Skill set: [’scoop’, ’stir’, ’drop food’, ’pull bowl closer’, ’open dumbwaiter’, ’close dumbwaiter’, ’start dumbwaiter’, ’put bowl into dumbwaiter’, ’DONE’, ’grasp spoon’, ’put spoon back’, ’move to purple bowl’, ’move to red bowl’]
@@
- Iteration 2 to 5 demonstrates how to scoop from red bowl to purlple bowl once, and iteration 6 to 9 repeats it.
Example user prompt: Includes environmental information, task instructions, and feedback messages from the Skill Affordance module, along with the current observation.
@@
Your task: Skill set: [’A. scoop’, ’B. stir’, ’C. drop food’, ’D. pull bowl closer’, ’E. open dumbwaiter’, ’F. close dumbwaiter’, ’G. start dumbwaiter’, ’H. put bowl into dumbwaiter’, ’I. DONE’, ’J. grasp spoon’, ’K. put spoon back’, ’L. move to white bowl’, ’M. move to blue bowl’]